OpenStudy (anonymous):

What is the solution of the equation (3x-4)/5=(2x+1)/10? Use Cross multiplication and show all work.

OpenStudy (anonymous):

\[\frac{3x-4}{5}=\frac{2x+1}{10}\]

OpenStudy (anonymous):

\[10(3x-4)=5(2x+1)\]

OpenStudy (anonymous):

\[30x-40=10x+5\] \[20x=45\] \[x=\frac{45}{20}=\frac{9}{4}\]